OpenStudy (daltonw):

What are the factors of 4x2 + 23x - 6? (4x - 6)(x + 1) (4x - 1)(x + 6) (4x - 2)(x + 3) (4x - 3)(x + 2)

OpenStudy (daltonw):

@Crissy15

OpenStudy (crissy15):

4x² + 23x - 6 = 4x² + 24x - x - 6 = 4x(x + 6) - 1(x + 6) = (4x - 1)(x + 6)

OpenStudy (daltonw):

thank you!!! i like that you wrote the proces out!! i have a lot more and this should help thanks!!!